OpenStudy (anonymous):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below. f(x) = 72(1.25)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer.

OpenStudy (anonymous):

please help I don't understand

OpenStudy (anonymous):

@zepdrix @thomaster @Luigi0210 @ParthKohli

zepdrix (zepdrix):

Is the function supposed to be:\[\Large\rm f(x)=72(1.25)^x\]yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Imagine we had a function like this:\[\Large\rm g(x)=72(1)^x\]Every year, the amount 72 would be unchanged, yes? It's not increasing or decreasing. The 1 represents 100%. It's staying at 100% all the time. A value larger than 1 will represent increase.\[\Large\rm f(x)=72(1.25)^x\]This 72 is growing to 125% of it's value each year. It's increasing by 25% each year.

OpenStudy (anonymous):

o thank u that makes a lot more sense